OpenStudy (anonymous):

Prove the identity (sinh x)^2 = (cosh 2x - 1) / (2)

OpenStudy (goformit100):

is 'h' showing the multiple angle ?

OpenStudy (anonymous):

No it is hyperbolic sin x

OpenStudy (anonymous):

pretty sure it is 100% algebra using \[\sinh(x)=\frac{e^x-e^{-x}}{2}\]

OpenStudy (anonymous):

Yes, but have a look at this solution, I don't know where the 1/2 comes from?

OpenStudy (larseighner):

The half is factored out. It "comes" from the 4 in the denominator which is then reduced to two.

OpenStudy (anonymous):

Ah, makes sense, but what about the -1? because -2(e^0) = -2 isn't it?

OpenStudy (larseighner):

It does equal -2. But (-2/2) = 1. The half came out of the denominator. It did not touch the numerator. All of the numerator has the common denominator of 2. So you pull -2/2 out with simple addition. (ething -2 + ething)/2 = (ething + ething)/2 - 2/2 = (ething + ething)/2 -1.

OpenStudy (anonymous):

Ohh, now I get it, thank you so much for the help :)